OpenStudy (anonymous):

Solve for x. 2x2 - 4x = 0 A. 0, -4 B. 0, -2 C. 0, 2 D. 2, 4

OpenStudy (anonymous):

C. 0, 2

OpenStudy (abb0t):

1. factor out a 2 and an "x" 2. use your tools of algebra to solve for one of the x's you have \(2x(x-2)\) HINT: one of your x values is zero.

OpenStudy (abb0t):

next time, @Brunette try to help them get the answer rather than just giving it to them. It helps them learn better.

OpenStudy (anonymous):

im sure they just want the answer

OpenStudy (abb0t):

Doesn't mean you have to give it to them :3

OpenStudy (anonymous):

first factor..... (2x-2) (x+2) ----- 2 = (x -0) zero out ...... x=0, x=-2

OpenStudy (anonymous):

just trying to help them out, i get that it can be really stressful by the end of the semester